In a closed economy, the MPC is 0.50. Investment changes by 100. The change in equilibrium GDP is
Answer #1

The given MPC is 0.50, investment changes by 100. So, the formula for the multiplier is:

=1/(1-MPC)

=1/(1-0.50)

= 2

The change in equilibrium GDP is = 2 * 100 =200